| Operation Summary for January 13, 2012 - January 16, 2012 |
|
Monday January 16 The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with the NuMI and MiniBooNE experiments taking beam. During the day shift, Operators reported that a MI RF cavity (MIRF9) had a very slow LCW leak. At 9:22 AM, a Linac RF station (LRF2) tripped off; it reset. Operations established beam to the Meson FTBF experiment T-992 at 10:07 AM. At 11:05 AM, LRF4 tripped off; it reset. Operators reported at 3:25 PM that the H- Source was failing. Throughout the shift Operators tuned the 400 MeV line and Booster to keep up with the source fluctuations. During the evening shift, Operations terminated T-992 beam at 9:55 PM. Operators reported at 11:53 PM that the H- Source had stabilized a bit. Tuesday January 17 The midnight shift began with Operations sending beam to the NuMI and MiniBooNE. During the midnight shift, at 2:40 AM, LRF5 began tripping off and being reset repeatedly until 3:08 AM when Operators reported hearing a loud pop; Operators called in a Linac technician. At 4:55 AM, the Linac tech reported that he had changed the LRF5’s Thyratron, but there were still problems with the station that held off Linac beam. During the day shift, at 8 AM, Neutron Therapy Facility personnel began preparing for treating a patient later this morning. A Linac person reported at 8:03 AM that they were replacing the power amplifier for LRF5. Operators prepared the Muon Rings for access at 8:17 AM. A Recycler kicker magnet (R:KPS3A) tripped off on a magnet temperature low Fluorinert alarm; MI personnel reported that the Fluorinert would be filled during the access on Wednesday. Operations reported that LRF5 was ready to operate at 11:37 AM, but problems with BRF16 held off beam. A Booster RF person reported that BRF16 had a bad modulator power supply; they bypassed the station and began repairs. Operations established beam to MiniBooNE at 11:56 AM and to NuMI at 12:02 PM. Operations established beam to T-992 at 12:09 PM. At 12:43 PM, LRF5 tripped off; it reset. At 1:47 PM, BRF4 tripped off; it reset. During the evening shift, at 4:31 PM, LRF5 tripped off; it reset. A MI person completed a study at 5:15 PM that he had started during the day shift. At 6:16 PM, LRF4 tripped off for the first of two times in four hours; it reset. Operations disabled T-992 beam at 8:49 PM; this completed T-992’s present run at Fermilab. Wednesday January 18 The midnight shift began with Operations sending beam to the NuMI and MiniBooNE. During the midnight shift, at 1:42 AM, a Booster vacuum pump (B:IPRF15) tripped off and brought down two RF stations (BRF3 and 4) with it; everything reset. At 2 AM, a Linac Klystron RF station (KRF1) tripped off; it reset. At 2:29 AM, LRF2 tripped off; it reset. Operators reported at 5:16 AM that a Linac quadrupole power supply’s (L:QPS216) output would drift high for about 30 seconds and then return to normal; Operators lowered the quadrupoles setting just slightly. Operations turned off beam at 7 AM in preparation for the eight-hour Booster access. At 7:10 AM, Operators prepared MiniBooNE for access. Operators had Linac, Booster, MI, the Transfer Hall, and F-sector reading for access by 7:38 AM. Operations Coordinator (OpsCo) Plans for Today The OpsCo wants everyone involved in today’s access to work safely. The access ends at 4 PM. |
|
More Information Comments and Suggestions |
| last modified 01/18/2012 email Fermilab |
| Security, Privacy, Legal | Use of Cookies | |